• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name JP. Chen

    Log entry time 14:15:24 on July21,2003

    Entry number 105859

    Followups:

    keyword=Shift summary

    8:00 continuing kin. 5.12, BHWP out, target at 270deg

    - run 2622/22622 7:57-8:10 3M coda events total 18M
    2623/22623 8:13-8:30 3M coda events 21M
    2624/22624 8:30-8:43 3M coda events 24M
    2625/22625 8:44-8:58 3M coda events 27M
    2626/22626 8:59-9:14 3M coda events 30M
    2627/22627 9:18-9:25 1.3M coda events 31M

    9:30 requested to have BHWP in, taking data at kin 5.12, BHWP in, target 270.

    -run 2628/22628 9:30-9:44 3M coda events total 3M
    2629/22629 9:47-10:02 3M coda events 6M
    2630/22630 10:03-10:10 1M coda event 7M

    10:10-12:00 beam study
    10:15 Target: transverse lasers off, moved to empty poistion, rotated from 270 to 0,
    longitudinal lasers on.
    11:00 Radcon survey for controlled access.
    11:15-11:24 NMR=41.5%,EPR=40.6%.
    Turned off longitudinal lasers, moved to 3He position, rotated to 270, turned on
    transverse lasers.
    Controlled access: 1)Vince to set spectro-analyzer for transverse laser wavelength check. 2) Bogdan to check? 3) Zafer from MCC to look at our beam line. 4) Rob to check S1m on left arm.
    12:25 Hall A back to beam permit
    12:10- Left DAQ crashed while doing a cosmic run to check S2m.
    12:30 asked for controlled access to check left DAQ
    12:35 noticed right DAQ runcontrol is gone. Restarted. OK.
    13:30 Bodo and Rob finished left DAQ work. Had trouble to close the detector hut door.
    14:15 With Jack's help, door was closed. back to beam permit